OpenStudy (anonymous):

How come you are recommended to find the angle opposite to the largest side when using the cosine law in a SideSideSide Triangle?

OpenStudy (lgbasallote):

dont forget to write the source rohangrr...

OpenStudy (dumbcow):

possibly because the angle used in Law of cosines is opposite of side being solved for

OpenStudy (anonymous):

It has something to do with obtuse and acute angles

OpenStudy (anonymous):

Theory: This triangle has internal angles 'A', 'B' and 'C', and sides of length 'a', 'b' and 'c':  If three of these six measurements are known, then it may be possible to find the other three. This is called 'solving' the triangle, and this topic will show you how to solve triangles for the three unknown angles when the three side lengths 'a', 'b' and 'c' are known. Some Formulas to solve: 1. The sum of the internal angles equals 180º ... A + B + C = 180º 2. The 'sine rule' ... a/sin A =b/sine B =c/sin C 3. The 'cosine rule' ... a² = b² + c² - 2bc cosA or b² = a² + c² - 2ac cosB or c² = b² + a² - 2ba cosC